# ADA-Track: End-to-End Multi-Camera 3D Multi-Object Tracking with Alternating Detection and Association

### v1

@@ Abstract
Many query-based approaches for 3D Multi-Object Tracking (MOT) adopt the tracking-by-attention paradigm, utilizing track queries for identity-consistent detection and object queries for identity-agnostic track spawning. Tracking-by-attention, however, entangles detection and tracking queries in one embedding for both the detection and tracking task, which is sub-optimal. Other approaches resemble the tracking-by-detection paradigm, detecting objects using decoupled track and detection queries followed by a subsequent association. These methods, however, do not leverage synergies between the detection and association task. Combining the strengths of both paradigms, we introduce ADA-Track, a novel end-to-end framework for 3D MOT from multi-view cameras. We introduce a learnable data association module based on edge-augmented cross-attention, leveraging appearance and geometric features. Furthermore, we integrate this association module into the decoder layer of a DETR-based 3D detector, enabling simultaneous DETR-like query-to-image cross-attention for detection and query-to-query cross-attention for data association. By stacking these decoder layers, queries are refined for the detection and association task alternately, effectively harnessing the task dependencies. We evaluate our method on the nuScenes dataset and demonstrate the advantage of our approach compared to the two previous paradigms. Code is available at https://github.com/dsx0511/ADA-Track.
@@ 1 Introduction
Figure 1: Paradigm comparison of query-based end-to-end MOT. Our proposed paradigm (1(c)) leverages the advantages of the coupled architecture of tracking-by-attention (1(a)) and the decoupled task-specific queries of tracking-by-detection (1(b)).
Accurate and consistent 3D Multi-Object Tracking (MOT) is critical for ensuring the reliability and safety of autonomous driving. Recently, vision-centric perception solely relying on multi-view cameras has garnered significant attention in the autonomous driving community, thanks to lower cost of sensors and the advancements of transformers for computer vision. Within this domain, two predominant approaches have emerged: one transforms multi-view features into an intermediate dense Bird’s-Eye View (BEV) representation [35, 26, 64, 23, 31], while the other leverages object queries [8] that directly interact with the multi-view images [42, 61, 51, 58, 66] to construct an object-centric representation. Owing to the advantages in modeling object motion, the latter has been extended to query-based MOT in many works [42, 61, 51, 58, 66].
Among the query-based MOT approaches, the majority adopts the tracking-by-attention (TBA) paradigm [42, 61]. As illustrated in Figure 1(a), TBA utilizes track queries (colored squares) to consistently detect the same identity across frames and introduces object queries (white squares) to initialize tracks for newly appearing objects in each frame. However, this highly entangled design is sub-optimal for balancing the detection and tracking performance. Firstly, each track query consisting of a single embedding is tasked with accomplishing both, detection and tracking, while the two tasks share the same network architecture. Furthermore, track queries for identity-aware tracking and object queries for identity-agnostic detection are also processed by identical network weights. We argue that such an approach is sub-optimal for extracting task-specific information from the single query representation. Secondly, data association is implicitly addressed using self-attention between all queries. Although it effectively integrates information of query relations into query refinement, a notable drawback arises during inference. The network only outputs one confidence score for each object, but it is unclear whether it represents a detection or an association confidence. This requires sophisticated manually tuned post-processing.
@@
In this paper, we argue that detection and tracking is a chicken-and-egg problem: accurate detection enables robust initialization and straightforward association to tracks, while well-established tracks incorporate temporal context to mitigate potential detection errors. Our method elegantly addresses this challenge by leveraging synergies in both tasks while decoupling them. We propose ADA-Track, a novel query-based end-to-end multi-camera 3D MOT framework that conducts object detection and explicit association in an alternating manner, as shown in Figure 1(c). We propagate track queries across frames representing a unique object instance, while generating decoupled detection queries that detect all objects in each frame. Inspired by [16], we propose a learned data association module based on an edge-augmented cross-attention [27]. In this module, edge features between track and detection queries represent association information. These features are incorporated into attention calculations, updated layer-by-layer, and further used to output affinity scores. Different from [16], we include appearance features in the nodes and geometric features in the edges, resulting in a fully differentiable appearance-geometric reasoning. We then integrate the learned association module into each transformer decoder layer of a query-based multi-camera 3D detector, e.g. DETR3D [55]. In this way, the decoder layer sequentially conducts a query-to-image cross-attention to refine query representations for object detection and a query-to-query cross-attention to refine query and edge representations for data association. By stacking the decoder layers, iteratively refined query and edge features provide useful information to each other, resulting in a harmonized optimization of the detection and tracking task.
@@
We evaluate our method on the nuScenes dataset [6] and compare our proposed alternating detection and association paradigm with approaches based on the other two paradigms. While achieving state-of-the-art performance, our proposed paradigm can easily be combined with various query-based 3D detectors.
@@ 2 Related Work
Multi-camera 3D detection
Current research on multi-camera 3D object detection falls into two major categories. The first category transforms multi-view image features into a dense Bird’s-Eye View (BEV) representation using CNNs [48, 26, 25, 33] or transformers [35, 59]. Although it has been demonstrated that temporal fusion in BEV effectively boosts the detection performance [25, 35, 47] or supports downstream tasks [23, 64, 31], such a structured representation may struggle to effectively model object motion. The second category contains works that follow DETR [8], where sparse object queries interact with multi-view images [55, 40, 17]. This sparse query-based representation facilitates effective object-centric temporal fusion by interacting queries with multi-frame sensor data [38] or propagating queries across frames [54, 39]. Consequently, end-to-end detection and tracking methods built on top of query-based detectors have emerged as a popular choice.
Tracking-by-attention
@@ 3 Approach
An overview of ADA-Track is shown in Figure 2. For each frame $t$, feature maps $F_{c}^{t}$ are extracted from multi-view images $I_{c}^{t}$ using a CNN for each camera $c$. A set of track queries $Q_{\text{T}}^{t}$ (depicted as colored squares in Figure 2) are propagated from the previous frame in order to consistently detect the same identity across frames. A fixed number of $N_{D}$ detection queries $Q_{\text{D}}^{t}$ (white squares in Figure 2) is randomly initialized and responsible to detect all objects in the current frame. Following recent works, we assign a 3D reference to each of those queries [55, 40]. The transformer decoder layer first conducts a self-attention between queries and an image-to-query cross-attention (e.g. DETR3D [55] or PETR [40]) to refine queries for the object detection task. Subsequently, a query-to-query edge-augmented cross-attention integrates both query types and edge features and refines them for the data association task. The overall transformer decoder layer is repeated $L_{d}$ times to alternately refine query and edge features for both detection and association task. Finally, a track update module associates the track and detection queries and generates track queries $Q_{\text{T}}^{t+1}$ for the next frame.
@@ 3.1 Joint detection and association decoder layer
Query-to-query self-attention
@@
Image-to-query cross-attention
@@
Query-to-query edge-augmented cross-attention
@@ 3.2 Track Update
Data association
We use different association schemes for training and inference. During training, we first match tracks and detections to ground-truth objects (details in Section 3.3). If a track and a detection query are matched to the same ground-truth identity, both queries are considered as a matched pair. All track queries that are unmatched with a ground-truth are terminated, while all detection queries that are matched to the newly appearing ground-truth objects spawn a new track. During inference, we apply an MLP followed by a sigmoid function on the final edge features $E^{(L_{d})}$ to estimate the affinity scores $S$ between all track-detection pairs, i.e. $S=\text{sigmoid}(\text{MLP}(E^{(L_{d})}))\in\mathbb{R}^{N_{D}\times N_{T}}$. Then, the score matrix $S$ is used as matching costs of a Hungarian Algorithm [28] to obtain a one-to-one matching. We heuristically keep the unmatched tracks for $T_{d}=5$ frames and mark them as temporally inactive tracks before termination. An unmatched detection box initializes a new track, if its confidence is higher than $\tau_{\text{new}}=0.4$.
Figure 3: Target assignments: Tracking-by-attention applies identity-guided matching for track queries and then matches detection queries to remaining ground truths using the Hungarian Algorithm. Our method employs the same matching rules for both query types, but detection queries are matched to all ground-truths.
Feature and box update
@@ 3.3 Training
Target assignment
Tracking-by-attention approaches use identity-guided matching for the track queries, while the Hungarian Algorithm [28] matches the remaining ground-truth boxes and the object queries. In this work, track and detection queries detect objects independently and are explicitly associated with each other, which requires isolated matching for both query types. To achieve that, we match all ground-truth identities with detection queries using the Hungarian Algorithm, in addition to the same identity-guided matching for track queries. Figure 3 illustrates this difference in target assignment. These matching results assign the targets for the box losses. In addition, we apply an explicit loss function for the association module, where we regard it as a binary classification problem. If a ground-truth identity is matched with both, a track query and a detection query, the edge between these two queries should be classified as positive. In all other cases, e.g. one of both queries is matched with $\emptyset$, or both queries are matched to different ground truths, the association target for this detection-track pair should be negative.
Loss function
For the bounding box loss in both query types, we follow [55, 40] and use the Focal Loss [37] as the classification loss $\mathcal{L}_{\text{cls}}$ and $\ell_{1}$-loss as the regression loss $\mathcal{L}_{\text{reg}}$. The track and detection queries contribute to separate loss terms. This results in $\mathcal{L}_{\text{cls,T}}$ and $\mathcal{L}_{\text{reg,T}}$ for the track and $\mathcal{L}_{\text{cls,D}}$ and $\mathcal{L}_{\text{reg,D}}$ for the detection part. For association, we use Focal Loss with $\alpha=0.5$ and $\gamma=1.0$, denoted as $\mathcal{L}_{\text{asso}}$. We include auxiliary losses after each intermediate decoder layer for all the previously mentioned loss terms.
@@
$\sum_{t=2}^{T}(\lambda_{\text{cls}}\mathcal{L}_{\text{cls,T}}^{t}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg,T}}^{t}+\lambda_{\text{loss}}\mathcal{L}_{\text{asso}}^{t}).$
@@
We use $\lambda_{\text{cls}}=2.0,\lambda_{\text{reg}}=0.25$ following existing works [55, 40] and $\lambda_{\text{asso}}=10.0$. The impact of the loss weight $\lambda_{\text{asso}}$ is further analyzed in the supplementary.

### v2

# ADA-Track++: End-to-End Multi-Camera 3D Multi-Object Tracking with Alternating Detection and Association

@@ Abstract
Many query-based approaches for 3D Multi-Object Tracking (MOT) adopt the tracking-by-attention paradigm, utilizing track queries for identity-consistent detection and object queries for identity-agnostic track spawning. Tracking-by-attention, however, entangles detection and tracking queries in one embedding for both the detection and tracking task, which is sub-optimal. Other approaches resemble the tracking-by-detection paradigm and detect objects using decoupled track and detection queries followed by a subsequent association. These methods, however, do not leverage synergies between the detection and association task. Combining the strengths of both paradigms, we introduce ADA-Track++, a novel end-to-end framework for 3D MOT from multi-view cameras. We introduce a learnable data association module based on edge-augmented cross-attention, leveraging appearance and geometric features. We also propose an auxiliary token in this attention-based association module, which helps mitigate disproportionately high attention to incorrect association targets caused by attention normalization. Furthermore, we integrate this association module into the decoder layer of a DETR-based 3D detector, enabling simultaneous DETR-like query-to-image cross-attention for detection and query-to-query cross-attention for data association. By stacking these decoder layers, queries are refined for the detection and association task alternately, effectively harnessing the task dependencies. We evaluate our method on the nuScenes dataset and demonstrate the advantage of our approach compared to the two previous paradigms.
Index Terms:
3D Multi-object tracking, multi-camera 3D perception, autonomous driving
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Accurate and consistent 3D Multi-Object Tracking (MOT) is critical for ensuring the reliability and safety of autonomous driving. Recently, vision-centric perception solely relying on multi-view cameras has garnered significant attention in the autonomous driving community, thanks to lower cost of sensors and the advancements of transformers for computer vision. Within this domain, two predominant approaches have emerged: one transforms multi-view features into an intermediate dense Bird’s-Eye View (BEV) representation [1, 2, 3, 4, 5], while the other leverages object queries [6] that directly interact with the multi-view images [7, 8, 9, 10] to construct an object-centric representation. Owing to the advantages in modeling object motion, the latter has been extended to query-based MOT in many works [11, 12, 13, 14, 15].
@@
Figure 1: Different paradigms of query-based MOT. Our proposed paradigm (1(c)) leverages the advantages of the coupled architecture of tracking-by-attention (1(a)) and the decoupled task-specific queries of tracking-by-detection (1(b)).
Among the query-based MOT approaches, the majority adopts the tracking-by-attention (TBA) paradigm, e.g. [11, 12]. As illustrated in Figure 1(a), TBA utilizes track queries (colored squares) to consistently detect the same identity across frames and introduces object queries (white squares) to initialize tracks for newly appearing objects in each frame. However, this highly entangled design is sub-optimal for balancing the detection and tracking performance. First, each track query consisting of a single embedding is tasked with accomplishing both, detection and tracking, while the two tasks share the same network architecture. Furthermore, track queries for identity-aware tracking and object queries for identity-agnostic detection are also processed by identical network weights. We argue that such an approach is sub-optimal for extracting task-specific information from the single query representation. Second, data association is implicitly addressed using self-attention between all queries. Although it effectively integrates information of query relations into query refinement, a notable drawback arises during inference. The network only outputs one confidence score for each object, but it is unclear whether it represents a detection or an association confidence. This requires sophisticated manually tuned post-processing.
@@
In this work, we argue that detection and tracking is a chicken-and-egg problem: accurate detection enables robust initialization and straightforward association to tracks, while well-established tracks incorporate temporal context to mitigate potential detection errors. Our method elegantly addresses this challenge by leveraging synergies in both tasks while decoupling them. We propose ADA-Track, a novel query-based end-to-end multi-camera 3D MOT framework that conducts object detection and explicit association in an alternating manner, as shown in Figure 1(c). We propagate track queries across frames representing a unique object instance, while generating decoupled detection queries that detect all objects in each frame. Inspired by [18], we propose a learned data association module based on an edge-augmented cross-attention [19]. In this module, edge features between track and detection queries represent association information. These features are incorporated into attention calculations, updated layer-by-layer, and further used to output affinity scores. Different from [18], we include appearance features in the nodes and geometric features in the edges, resulting in a fully differentiable appearance-geometric reasoning. We then integrate the learned association module into each transformer decoder layer of a query-based multi-camera 3D detector, e.g. DETR3D [7]. In this way, the decoder layer sequentially conducts a query-to-image cross-attention to refine query representations for object detection and a query-to-query cross-attention to refine query and edge representations for data association. By stacking the decoder layers, iteratively refined query and edge features provide useful information to each other, resulting in a harmonized optimization of the detection and tracking task.
A preliminary version of the proposed ADA-Track has been published in [20]. While our earlier work confirms the effectiveness of the novel tracking paradigm, we observe that the inherent properties of the softmax function in the cross-attention have some limitations for addressing the data association problem. In ADA-Track, we treat detections as queries and tracks as keys within the edge-augmented cross-attention. We follow standard attention mechanisms and apply softmax normalization across all existing tracks to compute the attention value towards a detection query, which in turn is used for affinity score estimation. However, such a detection query does not always have a corresponding track as an association target, e.g., due to newly appearing objects or occlusions. In these cases, applying softmax normalization gives a relatively high association score to some tracks, whereas the association score would be ideally the same for all tracks. This issue results in wrong associations and thus increases identity switches and degrades tracking performance. To address this issue, we propose ADA-Track++, an improved version of ADA-Track. ADA-Track++ introduces an auxiliary token as an additional association target for the detection queries in the edge-augmented cross-attention. This modification effectively mitigates the undesirable attention values produced by the softmax normalization. By incorporating the auxiliary token, the softmax normalization allocates some probability mass to this token, allowing the model to ignore irrelevant tracks by downgrading their attention. During training, we assign the auxiliary token as the association target for detection queries that should not be associated with any tracks.
In this improved version, we have also refined the original implementation [20], leading to improved performance and robustness of our method. To ensure the validity of our results, we have rerun all experiments and report the updated evaluation based on the refined implementation. In addition to these changes, we have conducted more comprehensive experiments to further analyze the proposed ADA-Track and its enhanced version, ADA-Track++, including a runtime and complexity analysis, reporting results for additional backbones, and providing additional ablation studies.
@@
We evaluate ADA-Track and ADA-Track++ on the nuScenes dataset [21]. We compare our proposed alternating detection and association paradigm with approaches based on the other two paradigms. While achieving promising performance, our proposed paradigm can easily be combined with various query-based 3D detectors. Compared to ADA-Track, ADA-Track++ achieves up to $0.7\%P$ and $1.1\%P$ higher AMOTA based on DETR3D and PETR respectively, demonstrating its stronger ability to handle data association.
@@ 2 Related Work
2.1 Multi-camera 3D detection
Current research on multi-camera 3D object detection falls into two major categories. The first category transforms multi-view image features into a dense Bird’s-Eye View (BEV) representation using CNNs [22, 2, 23, 24, 25, 26] or transformers [1, 27]. Although it has been demonstrated that temporal fusion in BEV effectively boosts the detection performance [23, 1, 28, 29] or supports downstream tasks [4, 3, 5], such a structured representation may struggle to effectively model object motion. The second category contains works that follow DETR [6], where sparse object queries interact with multi-view images [7, 8, 9, 30]. This sparse query-based representation facilitates effective object-centric temporal fusion by interacting queries with multi-frame sensor data [31] or propagating queries across frames [10, 32]. Consequently, end-to-end detection and tracking methods built on top of query-based detectors have emerged as a popular choice.
2.2 Tracking-by-attention
@@ 3 ADA-Track
In this section, we introduce the original version of ADA-Track [20]. An overview is shown in Figure 2. For each frame $t$, feature maps $F_{c}^{t}$ are extracted from multi-view images $I_{c}^{t}$ using a CNN for each camera $c$. A set of track queries $Q_{\text{T}}^{t}$ (depicted as colored squares in Figure 2) are propagated from the previous frame in order to consistently detect the same identity across frames. A fixed number of $N_{D}$ detection queries $Q_{\text{D}}^{t}$ (white squares in Figure 2) is randomly initialized and responsible to detect all objects in the current frame. Following recent works, we assign a 3D reference to each of those queries [7, 8]. The transformer decoder layer first conducts a self-attention between queries and an image-to-query cross-attention (e.g. DETR3D [7] or PETR [8]) to refine queries for the object detection task. Subsequently, a query-to-query edge-augmented cross-attention integrates both query types and edge features and refines them for the data association task. The overall transformer decoder layer is repeated $L_{d}$ times to alternately refine query and edge features for both detection and association task. Finally, a track update module associates the track and detection queries and generates track queries $Q_{\text{T}}^{t+1}$ for the next frame.
@@ 3.1 Joint detection and association decoder layer
3.1.1 Query-to-query self-attention
@@
3.1.2 Image-to-query cross-attention
@@
3.1.3 Query-to-query edge-augmented cross-attention
@@ 3.2 Track Update
3.2.1 Data association
We use different association schemes for training and inference. During training, we first match tracks and detections to ground-truth objects (details in Section 3.3). If a track and a detection query are matched to the same ground-truth identity, both queries are considered as a matched pair. All track queries that are unmatched with a ground-truth are terminated, while all detection queries that are matched to the newly appearing ground-truth objects spawn a new track. During inference, we apply an MLP on the final edge features $E^{(L_{d})}$ to estimate the affinity scores $S$ between all track-detection pairs, i.e. $S=\text{MLP}(E^{(L_{d})})\in\mathbb{R}^{N_{D}\times N_{T}}$. The score matrix $S$ is activated by a sigmoid function and used as matching costs of a Hungarian Algorithm [39] (HM), where the element with a pair is rejected in HM if its activated affinity score is lower than a threshold $\tau_{S}=0.3$. This produces a one-to-one matching, which produces matched track-detection pairs, unmatched tracks, and unmatched detections. We heuristically keep the unmatched tracks for $T_{d}=5$ frames and mark them as temporally inactive tracks before termination. An unmatched detection box initializes a new track if its confidence is higher than a threshold $\tau_{\text{new}}=0.4$, following the score threshold for spawning tracks from object queries in [35, 37].
Figure 3: Target assignments: Tracking-by-attention (a) applies identity-guided matching for track queries and then matches detection queries to remaining ground truths using the Hungarian Algorithm. Our method (b) employs the same matching rules for both query types, but detection queries are matched to all ground-truths.
3.2.2 Feature and box update
@@ 3.3 Training
3.3.1 Target assignment
As shown in Figure 3 (a), tracking-by-attention approaches use identity-guided matching for the track queries, where the bounding boxes of the same identity are always matched to the same track query, while the Hungarian Algorithm [39] matches the remaining ground-truth boxes and the object queries, similar to DETR [6]. In this work, track and detection queries detect objects independently and are explicitly associated with each other, which requires isolated matching for both query types. To achieve that, we use another matching scheme where the matching of both query types is independent, as shown in Figure 3 (b). We match all ground-truth identities with detection queries using the Hungarian Algorithm [39], in addition to the same identity-guided matching for track queries. These matching results define the targets for the bounding box losses.
In addition to the bounding box losses, we apply an explicit loss function for the association module, where we regard it as a binary classification problem. If a ground-truth identity is matched with both, a track query and a detection query, the edge between these two queries should be classified as positive. In all other cases, e.g. one of both queries is matched with $\emptyset$, or both queries are matched to different ground truths, the association target for this detection-track pair should be negative.
3.3.2 Loss function
For the bounding box loss in both query types, we follow [7, 8] and use the Focal Loss [64] as the classification loss $\mathcal{L}_{\text{cls}}=\mathcal{L}_{\text{FL}}$ and $\ell_{1}$-loss as the regression loss $\mathcal{L}_{\text{reg}}=\mathcal{L}_{\ell_{1}}$. The track and detection queries contribute to separate loss terms. This results in $\mathcal{L}_{\text{cls,T}}$ and $\mathcal{L}_{\text{reg,T}}$ for the track and $\mathcal{L}_{\text{cls,D}}$ and $\mathcal{L}_{\text{reg,D}}$ for the detection part. For association, we use Focal Loss with $\alpha=0.5$ and $\gamma=1.0$, denoted as $\mathcal{L}_{\text{asso}}=\mathcal{L}_{\text{FL}}$. We include auxiliary losses after each intermediate decoder layer for all the previously mentioned loss terms.
@@ 3.3.2 Loss function
$\sum_{t=2}^{T}(\lambda_{\text{cls}}\mathcal{L}_{\text{cls,T}}^{t}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg,T}}^{t}+\lambda_{\text{asso}}\mathcal{L}_{\text{asso}}^{t}).$
@@
We use $\lambda_{\text{cls}}=2.0,\lambda_{\text{reg}}=0.25$ following existing works [7, 8] and $\lambda_{\text{asso}}=10.0$. The impact of the loss weight $\lambda_{\text{asso}}$ is further analyzed in the experiments.
4 ADA-Track++
In this section, we present ADA-Track++, which improves the data association of ADA-Track by introducing an auxiliary token in the edge-augmented cross-attention.
4.1 Problem of Attention Normalization in Tracking
In a transformer’s attention layer, the softmax function is applied to normalize attention scores for each query, converting them into a probability distribution over all keys. These normalized attention scores are then used as weights to update the query features through a weighted sum of the values associated with the keys, where the normalization ensures the stability of this feature aggregation. As depicted in Figure 2, in the edge-augmented cross-attention of ADA-Track, we regard detections as queries that attend to tracks and aggregate features from them. Different from the standard cross-attention, the normalized attention weights are not only used to update the detection features but also to generate the affinity matrix $S$. In this setup, we aim for each detection query to primarily attend to the track query representing the same instance, while still slightly attending to other track queries to capture scene context or consider alternative association candidates. However, challenges arise when a detection query corresponds to a newly appearing object or a false positive. In such cases, no track should dominate the attention, but the normalized attention scores will still sum to one, potentially causing the detection query to focus on irrelevant tracks. To make things worse, even if all pre-softmax attention scores are low for certain detection queries, the highest score will be amplified after softmax (see Figure 4(a)). This results in an unreasonably large probability for assigning the detection to certain tracks, which leads to wrong assignments. The normalized attention can further mislead the feature aggregation and affinity estimation and negatively impact the overall performance.
(a) Attentions without auxiliary token.
(b) Attentions with auxiliary token (white).
Figure 4: The difference in attention before and after softmax normalization. Ground-truth track IDs are represented by unique colors. Figure 4(a): Although the yellow detection query does not have a corresponding track and its pre-softmax attentions are low everywhere, it still receives relatively high attention to the blue track after softmax normalization, which might disturb the correct association (blue track and blue detection). Figure 4(b): After introducing the auxiliary token, the normalized attention becomes much more reasonable.
4.2 Auxiliary token
To address the issue of normalized attention in the context of learning data association, we propose a simple yet effective solution. We introduce a randomly initialized auxiliary token $q_{\emptyset}\in\mathbb{R}^{d_{k}}$ with learnable weights at the first time stamp and propagate it across frames, similar to the track queries. In each decoder layer, the auxiliary token participates in the query-to-query self-attention while not attending to the images in the query-to-image cross-attention. In the edge-augmented cross-attention, we concatenate this auxiliary token to the tracks, i.e., the keys, representing a None identity and serving as an additional attention target for the detection queries. This produces an attention matrix $A\in\mathbb{R}^{N_{D}\times(N_{T}+1)}$, where $N_{D}$ and $N_{T}$ denote the number of detection queries and track queries, respectively. The inclusion of the auxiliary token allows the model to capture detection queries that should not be associated with any existing track queries. This can be depicted in Figure 4(b), where introducing the auxiliary token leads to a desirable attention distribution, which mitigates the risk of irrelevant track queries receiving disproportionate attention, thereby improving the overall data association process.
To ensure that the attention on the auxiliary token is highest when needed, we assign the auxiliary token as the association target in those cases. Since the association score $S\in\mathbb{R}^{N_{D}\times(N_{T}+1)}$ is derived from the attention matrix $A$, the supervision signal on $S$ will be back-propagated to the attention $A$. This forces detection queries with no association targets to attend to the auxiliary token. In ADA-Track, i.e., without the auxiliary token, the association target for each detection query could either be a one-hot vector, indicating a match with one track, or an all-zero vector, indicating no match with any track. With the inclusion of the auxiliary token in ADA-Track++, all association targets are represented as one-hot vectors. Consequently, we can employ categorical cross-entropy as an additional association loss term, which enforces one-hot affinities across tracks for each detection query. Given $Y_{ij}$ as the association target between the $j$-th detection and the $i$-th track, and $S_{ij}$ representing the affinity before any activation, the additional loss term can be formulated as
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N_{D}}\sum_{j=1}^{N_{T}+1}Y_{ij}\log\frac{\exp^{S_{ij}}}{\sum_{k=1}^{N_{T}+1}\exp^{S_{ik}}}.$ (4)
Then the overall association loss $\mathcal{L}_{\text{asso}}$ in Equ. 3 becomes a weighted sum of the existing Focal Loss and the cross-entropy loss in Equ. 4: $\mathcal{L}_{\text{asso}}=\mathcal{L}_{\text{FL}}+\lambda_{CE}\mathcal{L}_{\text{CE}}$, where $\lambda_{CE}=0.1$ is the weight controlling the contribution of the cross-entropy loss term $\mathcal{L}_{\text{CE}}$ to the overall association loss.